OpenStudy (anonymous):

A relative maximum of the function f(x)=(lnx)^2/x occurs at... Answer: e^2 Please show work and explain how to come to the answer.

OpenStudy (anonymous):

take the derivative, set it equal to zero and solve

OpenStudy (anonymous):

you got \(f'(x)\) ?

OpenStudy (anonymous):

I tried to do that. Would it be easier to use the product rule or the quotient rule?

OpenStudy (anonymous):

you have no choice, you have to use the quotient rule

OpenStudy (anonymous):

ok I'll try that.

OpenStudy (anonymous):

not too hard \[\left(\frac{f}{g}\right)'=\frac{gf'-fg'}{g^2}\] with \[f(x)=(\ln(x))^2, f'(x)=\frac{2\ln(x)}{x},g(x)=x, g'(x)=1\]

OpenStudy (anonymous):

Thanks I got it. :)